OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 − 4x + y2 + 8y = −4

OpenStudy (anonymous):

Yes, but I don't know how to do that.

OpenStudy (campbell_st):

do you know how to complete the square....?

OpenStudy (anonymous):

The general equation of a circle is:\[x^2+y^2+2gx+2fy+c=0\]

OpenStudy (anonymous):

where the center is:\[(-g,-f)\]

OpenStudy (anonymous):

so if you subtract 4 from both sides of the equation you get:\[x^2+y^2-4x+8y+4=0\]

OpenStudy (anonymous):

in your case 2g is -4 and 2f is 8 so the center is (-g,-f) which is (2,-4)

OpenStudy (anonymous):

so the center is (2,-4)

OpenStudy (anonymous):

now the radius is given by:\[\sqrt{g^2+f^2-c}\]

OpenStudy (anonymous):

ok @Hero

OpenStudy (anonymous):

ok so in your case g is -2 and f is 4 and c is 4. so the radius is:\[\sqrt{(-2)^2+4^2-4}\] that simplifies to 4 so the radius of the circle is 4

OpenStudy (anonymous):

there is also another equation of a circle which is:\[(x-h)^2+(y-k)^2=r^2\] where the radius is r and the center is (h,k) but in your problem it was in general form so I used the other circle equation